OpenStudy (anonymous):

PLEASE HELP!!!! FAST I need the vertical + horizontal asymptotes for f(x)=2x^2+7x-15/5x+x^2

OpenStudy (anonymous):

HA y=2 and va x-0 and x=5

jimthompson5910 (jim_thompson5910):

H.A.: y = 2, correct V.A. x = 0, correct V.A. x = -5 ... notice how it's negative

jimthompson5910 (jim_thompson5910):

x^2 + 5x = 0 x(x+5) = 0 x = 0 or x+5 = 0 x = 0 or x = -5 so that explains why it's x = -5 and not x = 5

OpenStudy (anonymous):

thank you so much!!! I wrote the wrong this on here haha

jimthompson5910 (jim_thompson5910):

wrote what wrong

OpenStudy (anonymous):

I need to check my work on my last problem. Would you be able to check it? I did it a million times today and took me hours/

jimthompson5910 (jim_thompson5910):

sure go for it

OpenStudy (anonymous):

a-f

OpenStudy (anonymous):

Thank you so much... Literally working on it for hours

jimthompson5910 (jim_thompson5910):

ok post what you did

OpenStudy (anonymous):

I got a cp of 6.1 about and concavities of all positive and increasing after 6.1 and decreasing before 6.1.

OpenStudy (anonymous):

aan IP at 4

OpenStudy (anonymous):

second derivative is 12x^2(x-4)

OpenStudy (anonymous):

I couldn't synthetically divide it

OpenStudy (anonymous):

stressing me out

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

f(x) = x^4-8x^3-16x+5 f ' (x) = 4x^3-24x^2-16 f '' (x) = 12x^2-48x

jimthompson5910 (jim_thompson5910):

The critical point is roughly 6.10724315175795

jimthompson5910 (jim_thompson5910):

you get this from solving 4x^3-24x^2-16 = 0 for x

jimthompson5910 (jim_thompson5910):

f(x) is increasing on the interval (6.10724315175795, infinity) this is a rough approximation of course

jimthompson5910 (jim_thompson5910):

f(x) is decreasing on the interval (-infinity, 6.10724315175795) and both of these facts can be seen from a graph or by using the first derivative test

OpenStudy (anonymous):

wait..when I drew the graph it looked concave up

jimthompson5910 (jim_thompson5910):

it is concave up

jimthompson5910 (jim_thompson5910):

well for the most part it is

OpenStudy (anonymous):

oh lol the negative infinity one confused me

jimthompson5910 (jim_thompson5910):

actually it's concave down on the interval (0,4)

OpenStudy (anonymous):

it passes through 0,$?

jimthompson5910 (jim_thompson5910):

12x^2-48x = 0 12x(x-4) = 0 12x = 0 or x-4 = 0 x = 0 or x = 4

OpenStudy (anonymous):

0,4

jimthompson5910 (jim_thompson5910):

the possible inflection points are at x = 0 and x = 4

OpenStudy (anonymous):

but only 4 works right

OpenStudy (anonymous):

where is it concave up and down between what intervals

jimthompson5910 (jim_thompson5910):

Now see if there is a sign change as you move through x = 0 f '' (x) = 12x^2-48x f '' (-1) = 12(-1)^2-48(-1) f '' (-1) = 60 f '' (x) = 12x^2-48x f '' (1) = 12(1)^2-48(1) f '' (1) = -36 this means that x = 0 is definitely an inflection point

OpenStudy (anonymous):

what are the concavity intervals

jimthompson5910 (jim_thompson5910):

notice how f '' (-1) = 60 is positive, so it's concave up on (-infinity, 0)

jimthompson5910 (jim_thompson5910):

f '' (1) = -36 is negative, so it's concave down on (0,4)

jimthompson5910 (jim_thompson5910):

now plug in a value to the right of x = 4 f '' (x) = 12x^2-48x f '' (5) = 12(5)^2-48(5) f '' (5) = 60 f '' is once again positive, so this confirms the other point of inflection is at x = 4 and it's concave up on (4, infinity)

OpenStudy (anonymous):

the confusing part was the 0,4

jimthompson5910 (jim_thompson5910):

that's the interval where f is concave down

OpenStudy (anonymous):

is the graph rising from 6.1,infinity

jimthompson5910 (jim_thompson5910):

yeah it's increasing on this interval

OpenStudy (anonymous):

and decreasing from -infinity,6.1

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

would you be able to draw the graph for me lol. I have ridiculous numbers

OpenStudy (anonymous):

I plugged 6.1 in and got like -315

OpenStudy (anonymous):

I need the graph and the range and I'll be ok

jimthompson5910 (jim_thompson5910):

posted the graph

jimthompson5910 (jim_thompson5910):

that's the graph of f(x)

jimthompson5910 (jim_thompson5910):

the range will be the set of y values that are larger than the smallest possible y value (ie the absolute min)

OpenStudy (anonymous):

can you help me lable this.. my teacher wants certain points on here and I didn't understand how he got them

OpenStudy (anonymous):

I am taking a month long course it is so accelerated for me

jimthompson5910 (jim_thompson5910):

well you got the critical point already right?

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

the critical point for f(x)

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the other two points are the inflection points

jimthompson5910 (jim_thompson5910):

to find them, you plug in x = 0 and x = 4 into f(x)

OpenStudy (anonymous):

I got 0,-5 but it is not on the graph

jimthompson5910 (jim_thompson5910):

it should be (0,5)

jimthompson5910 (jim_thompson5910):

not sure how you got -5

OpenStudy (anonymous):

I mean positive 5

OpenStudy (anonymous):

I plugged 6.1 in and got neg 523

jimthompson5910 (jim_thompson5910):

-523.8639, so close enough I guess

OpenStudy (anonymous):

the inflection points I am confused with

jimthompson5910 (jim_thompson5910):

well you found one and that was (0,5)

jimthompson5910 (jim_thompson5910):

the other is found when you plug in x = 4

OpenStudy (anonymous):

is it ok if it is not on my graph?

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

4,-315

OpenStudy (anonymous):

and the range?

OpenStudy (anonymous):

thanks for the help

jimthompson5910 (jim_thompson5910):

plug in x = 6.1 to get the smallest possible y value (ie the absolute min) this will help find the range

OpenStudy (anonymous):

-524.8,infinity?

OpenStudy (anonymous):

523

jimthompson5910 (jim_thompson5910):

more like -523.8639 actually, but you get the idea

jimthompson5910 (jim_thompson5910):

so the range is [-523.8639, infinity)

jimthompson5910 (jim_thompson5910):

I guess -523.86 is more accurate

OpenStudy (anonymous):

thanks a bunch man. why did my teacher give me such a tough example?

OpenStudy (anonymous):

haha

jimthompson5910 (jim_thompson5910):

that's what some teachers do

OpenStudy (anonymous):

I really appreciate the patience.

jimthompson5910 (jim_thompson5910):

and that's a good thing sometimes, you want your teacher to make you work hard

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

should I put a place marker near 0,8?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

the graph appears to be going around there or should I ignor it

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (anonymous):

the right half of the graph crosses x=8 something in the drawing

OpenStudy (anonymous):

I might be over analyzing this

jimthompson5910 (jim_thompson5910):

what about it?

OpenStudy (anonymous):

should I put a point there or leave it alone

jimthompson5910 (jim_thompson5910):

hmm it doesn't explicitly say "find the roots of f(x)", so I would say no

jimthompson5910 (jim_thompson5910):

don't worry about it

OpenStudy (anonymous):

ok haha perfect. Thanks a bunch!!!